I don't know what it is called but i do not know how to do this: Example x-3=3x+5
Found 2 solutions by benazir.sj@gmail.com, rfer:
Answer by benazir.sj@gmail.com(109)

x-3=3x+5
here you have to solve for x
x-3=3x+5
add 3 on both sides
x-3+3=3x+5+3
x=3x+5
to isolate x subtract 3x on both sides
x-3x=3x-3x+8
-2x=8
-2x/-2=8/-2
x=-4

Answer by rfer(16322)

x-3=3x+5
-2x=8
x=-4